OpenStudy (anonymous):

A woman throws a baseball from the front of an airplane to the back. From the perspective of the woman, the baseball travels at an average speed of 15 m/s. To an observer on the ground, the airplane is traveling at a constant speed of 820 km/h. What is the speed of the baseball from the perspective of the ground observer?

OpenStudy (anonymous):

So for these questions, you add if they're going in the same direction and you subtract if it's in the opposite direction for the ground observer, since the ball is moving in the opposite direction of the plane, it would be 820 km/h - 15 m/s 15 m/s * 60 s/min * 60 min/hour = 54000 m/h, 54000 m/h * 1000 m/km = 54 km/h 820 km/h - 54 km/h = 768 km/h to the ground observer By the way, I don't think it's safe for someone to throw a baseball that fast in an airplane

OpenStudy (anonymous):

the choices are 228 m/s 243 m/s 213 m/s 840 m/s

OpenStudy (roadjester):

I have to agree with @greenleaf800073 's analysis. 768km/hr is an option

OpenStudy (anonymous):

how do you convert from Km/hr to m/s?

OpenStudy (roadjester):

There are 1000m in a km 60 min in an hour 60s in a minute.

OpenStudy (roadjester):

What did you get?

OpenStudy (anonymous):

228 i think idk haha

OpenStudy (roadjester):

...

OpenStudy (roadjester):

768 \(\dfrac {km}{hr}\) X \(\dfrac {1000m}{km}\) X \(\dfrac {1hr}{3600s}\)=?

OpenStudy (anonymous):

213.3

OpenStudy (roadjester):

\(\huge ☺\)

OpenStudy (anonymous):